# A Heuristic Algorithm Based on Beam Search and Iterated Local Search for the Maritime Inventory Routing Problem

### v1

@@ 4.2 Beam Search Algorithm
The Harpy Speech Recognition System (Lowerre, 1976) was the first to implement what later became known as Beam Search (Ow and Morton, 1988), introducing a heuristic-based approach to efficiently explore and prune the search space in speech recognition tasks. Although the procedure was originally referred to as the Locus Model of Search, the term Beam Search was already used by Reddy (1977). BS is a method for exploring decision trees, especially when dealing with a large solution space. It involves methodically expanding a limited set of solutions simultaneously, aiming to increase the chance of finding a high quality solution with limited searching effort.
@@
During the Initialization stage, the algorithm establishes the starting parameters, including the initial node, which will always be an empty list, alongside the beam width. In the Node Expansion phase, successors are formed for each active node. Each successor is a partial solution. For a given node and its corresponding solution, all valid calls that can be appended to the solution are considered. Then, for every call, a new node is created, incorporating all previous calls along with the new one. Consequently, at the $i$-$th$ level of the generated tree, every solution consists of $i$ calls. In the Successor Evaluation phase, a greedy heuristic function is employed to assess the quality of each successor. Next, the Beam Selection phase involves selecting the most promising successors based on the heuristic evaluations while adhering to the beam width limit. The process continues with the Iteration stage, repeating until no further successors can be generated from the final level within the planning horizon.
@@
In the present study, the BS heuristic is configured to accept three distinct parameters as input: number of nodes allowed at a given level (N), maximum number of children per node (w) and the number of greedy solutions for each node (q).
The initial parameter to consider is $N$, which represents the number of nodes allowed within a specified level. The challenge lies in selecting a value that ensures an adequate array of potential initial solutions while keeping size within reasonable limits, which would otherwise result in higher CPU time. An increase in the number of ports and vessels, along with an extended instance time horizon, has been observed to lead to an increase in CPU time. For the test experiments, the $N$ used ranged according to the characteristics of each test. For the final experiments, different values of $N$ were evaluated based on the CPU time.
The second parameter, denoted as $w$, represents the upper limit on the number of children that can be generated by a single node. This parameter plays an important role in maintaining the diversity and heterogeneity among the solutions.
The third parameter, denoted as $q$, dictates the number of greedy solutions generated during the process, where one of them is deterministic, and $q-1$ is nondeterministic. This parameter directly interfaces with the greedy heuristic algorithm, which is used to assess the BS calculations. Specifically, the greedy heuristic is executed once following a deterministic approach, and subsequently, it is repeated $q-1$ times incorporating randomness into its execution to introduce variability in the solution space, as will be elaborated on in later sections.
@@
One problem already mentioned in Section 4.1 is the fact that several identical solutions are represented in different ways. A simple procedure was devised with the aim of reducing the occurrence of solutions that share the same cost evaluations, thereby improving the efficiency of BS, as it is undesirable to occupy unnecessary space in the BS level. This procedure functions by maintaining a dictionary where the keys correspond to the scores of potential solutions. By referencing this dictionary, it is possible to determine whether a particular score has been assigned to any solution. In such cases, these solutions are excluded from further consideration. An unlikely issue might arise with the presence of two distinct solutions that share the same cost, due to the intricate nature of the problem. Nevertheless, even if this occurs, the benefits continue to outweigh the disadvantages.
@@ 4.3 The Greedy Algorithm
To effectively compare the partial solutions, i.e., the successors produced during a BS execution, a greedy heuristic is used to construct a complete solution starting from a given partial solution, as presented in Algorithm 1. The greedy heuristic algorithm is built in both deterministic and nondeterministic versions. The heuristic constructs a feasible solution using a greedy method, making each decision by selecting the most favorable immediate choice according to a defined criterion. The criterion used involved choosing the next port at risk of violating inventory levels and determining which vessel could arrive there first. Therefore, the path was designed to prioritize, in the subsequent Call, the addition of the port at risk of inventory violation and the vessel capable of reaching this port in the earliest possible time.
@@
3while  $t\leq timehorizon$  do
4
5      $nextport\leftarrow$ next port to violate
6      $nextvessel\leftarrow$ vessel arrives first
7      $s_{partial}\leftarrow$ (nextport, nextvessel)
@@
To enhance flexibility and solution exploration, a nondeterministic version of the greedy algorithm is implemented, introducing randomness at two stages of the process: the port selection and the vessel selection. This randomness mechanism allows the algorithm to explore a wider range of possible paths, helping to escape suboptimal solutions and potentially improving the quality of the final solution.
In the nondeterministic version, instead of choosing the port that first violates its capacity constraint or the ship that first serves the selected port, the inverse of violation times is used as a probabilistic criterion. This means that ports with earlier violation times and vessels with earlier arrival times have a higher probability of being selected, although the choice is not fixed. This approach enables the algorithm to allow solutions that may seem less promising in the short term but, when combined, could lead to less infeasible and higher-quality solutions.
To incorporate this strategy into BS, the heuristic is executed $q$ times. In one execution, the process is entirely deterministic, following the original selection criterion. In the remaining $q-1$ executions, randomness is applied. At the end of these $q$ executions, the median of the solutions obtained is selected to compose the final score.
Notably, throughout the BS procedure, multiple complete solutions are assessed via this greedy heuristic. To ensure that valuable solutions are retained, the decision was made to preserve the top $N$ solutions identified by the greedy heuristic, in case they surpass the ultimate solution of the BS.
@@ 4.4 Local search
The local search procedure is based on a Randomized Variable Neighborhood Descent (RVND) scheme, where each neighborhood is explored in sequence until an improvement is found. If no improvement is found, the procedure moves on to the next neighborhood, following a random order. If any neighborhood results in a better solution, the process is restarted. Furthermore, our implementation uses the first improvement strategy.
An important consideration when considering these neighborhoods is that, in some cases, certain Calls may be displaced outside the planning horizon. When this happens, these Calls are ignored during the solution evaluation as they do not contribute to the cost or feasibility within the planned period.
Six neighborhood structures were implemented and they are described below.
@@
  • Remove: removes a Call along with the next Call of the same ship. This procedure is carried out to preserve vessel parity. For example, consider the following sequence of Calls: (P1, V1), (P2, V3), (P3, V1), (P1, V1), (P1, V2), (P1, V3), (P3, V2). If the Call (P3, V1) is removed, it becomes necessary to eliminate another occurrence of V1 to maintain the producer-consumer balance. In this case, both (P3, V1) and (P1, V1) are removed.
@@ 4.5 Iterated Local Search
ILS (Lourenço et al., 2003) applies an iterative process that alternates between perturbation and local search procedures. In each iteration, a perturbation is randomly applied to the current solution, which involves making a controlled change in its structure, by altering the routing structure of the current solution, triggering a change in the cost value of the solution. The perturbation mechanism selects one the six neighborhoods defined in Section 4.4 and applies a move at random. In our implementation, such a mechanism is called twice in a row. This perturbation generates a new candidate solution which is then refined using a local search to improve its quality within the neighborhood of the current solution. The algorithm then checks if the new solution will be accepted by a pre-established Simulated Annealing (Kirkpatrick et al., 1983) criterion. Once a solution is accepted, the algorithm updates the current solution and checks for improvement. If no improvement occurs, a counter increments. When this counter exceeds the allowed number of non-improving iterations, the solution is restored, and the counter resets. This process repeats until the maximum number of iterations is reached, at which point the final solution is returned.
@@ 5 Computational Experiments
The proposed algorithm was coded in Julia (version 1.10.5) and executed on an Intel(R) Core(TM) i7-8700K CPU @ 3.70GHz running Ubuntu 20.04, and using the GNU Parallel package (Tange, 2011) to distribute one task per thread.
@@ 5.1 Instance Details
Benchmark instances and their best-known solutions are maintained on the MIRPLib website (http://mirplib.scl.gatech.edu). The metadata includes high-level information for each instance. Most of the parameters are self-explanatory and were presented previously in Section 3. To reader is referred to Papageorgiou et al. (2014) for further explanation.
A software package has been created to access and use the public instances available in MIRPLib. Upon publication of the article, this package will be published for public use.
This paper addresses Group 2 instances, representing a single-product MIRP, considering only one port per region and never involving split pickups or split deliveries.
All instances have one loading region (LR) and different number of discharging regions (DR). The instance LR1DR02VC01V6a has one load region, two discharging regions, six vessels of one vessel class available, while instance LR1DR04VC03V15b has one load region, four discharging regions, 15 vessels of three vessel classes. The final letter “a” or “b” refers to the production/demand rates in the ports. If the letter is “a” the rates are constant, and if it is “b” the rates vary during the time horizon.
According to the classification presented by Papageorgiou et al. (2018), instances are grouped into three distinct categories based on their difficulty: Easy (E), Medium (M), and Hard (H). An instance is categorized as Hard (H) if, within a time frame of 1800 seconds, no commercial MILP solver operating with default parameter settings achieves a reduction greater than 10% of the gap, as defined in the equation 9. Conversely, an instance is considered Easy (E) if at least one commercially available MILP solver can reduce the gap by more than 90% within the same time limit using the default CPLEX settings. Instances that do not meet the criteria for the Easy or Hard categories are classified as Medium (M).
The main metric used to evaluate different solution methodologies involves computing the average proportion of the relative gap that is progressively closed as time elapses. This average is calculated for all instances classified as Easy, Medium, or Hard. In this context, the relative gap is quantitatively defined as
$\min\left\{\frac{z_{\text{method}}-z_{\text{best}}}{z_{\text{best}}},1\right\},$ (9)
where $z_{method}$ refers to the value of the objective function obtained from the method currently being assessed and $z_{best}$ represents the value of the objective function corresponding to the optimal solution identified so far in the literature.
@@ 5.2 Parameter Tuning
This section outlines how the parameters of the BS approach, the greedy algorithm, and ILS were calibrated.
@@ 5.2.1 BS Parameters
To determine a promising value for the maximum number of children per node, $w$, a set of experiments was performed. Table 1 reports the average and median objective value gaps obtained from tests in which the parameter $w$ was varied from 2 to 7. These percentages reflect the aggregated gap in objective values across all runs. The best performance was achieved with $w=2$, which resulted in the lowest average and median gaps. All tests were conducted on the Easy instances, with $N=100$ and 10 seeds per instance for each value of $w$.
Table 1: Gap analysis for different $w$ values
@@ 5.2.2 Greedy Heuristic Parameters
Tests were carried out to determine the value of the number of greedy solutions $q$ for each node, with the aim of maximizing performance relative to CPU time. To this end, the number of nodes allowed per level in the BS ($N$) was set to be inversely proportional to $q$, ensuring that the CPU time remained approximately constant. Figure 4 illustrates the results, showing the calibration of $q$ for the Easy instances. The score indicates the average gap obtained across 10 different seeds. The best results were achieved for $q=3$, as it yielded the smallest average gap.
@@ 5.2.3 ILS Parameters
To tune the ILS parameters, the HyperTuning.jl package (de Dios et al., 2021) was used for automated hyperparameter optimization in Julia. In order to optimize the ILS parameters, 15 instances together with five seeds were selected. A total of 120 trials were conducted and the parameter optimization process lasted for 72 hours. The parameter range, step, and results are detailed in Table 2. The results indicated the final set of parameter values that corresponds to the configuration that achieved the best overall performance.
@@
Parameter Definition range step result
$sa$-$in$ SA initial probability 0.75–0.99 $0.01$ $0.79$
$sa$-$fi$ SA final probability 0.01–0.10 $0.01$ $0.01$
$ils$-$\eta$ ILS number of iterations 10–1000 10 640
$ils$-$re$ ILS number of non-improving iterations before solution restore 1–20 1 4
$ils$-$pb$ ILS number of perturbations 1–5 1 2

### v2

@@ 4.3 Beam Search Algorithm
The Harpy Speech Recognition System (Lowerre, 1976) was the first to implement what later became known as Beam Search (Ow and Morton, 1988), introducing a heuristic-based approach to efficiently explore and prune the search space in speech recognition tasks. Although the procedure was originally referred to as the Locus Model of Search, the term Beam Search was already used by Reddy (1977). BS is a method for exploring decision trees, especially when dealing with a large solution space. It involves methodically expanding a limited set of solutions simultaneously, aiming to increase the chance of finding a high-quality solution with limited searching effort.
@@
During the Initialization stage, the algorithm establishes the starting parameters, including the initial node, which will always be an empty list, alongside the beam width. In the Node Expansion phase, successors are formed for each active node. Each successor is a partial solution. For a given node and its corresponding solution, all valid calls that can be appended to the solution are considered. Then, for every call, a new node is created that incorporates all previous calls together with the new one. Consequently, at the $i$-$th$ level of the generated tree, every solution consists of $i$ calls. In the Successor Evaluation phase, a greedy heuristic function is employed to assess the quality of each successor. Next, the Beam Selection phase involves selecting the most promising successors based on the heuristic evaluations while adhering to the beam width limit. The process continues with the Iteration stage, repeating until no further successors can be generated from the final level within the planning horizon. In the Termination and Output generation phase, the best solutions generated in all steps are selected.
@@
In the present study, the BS heuristic is configured to accept three different parameters as input, which have an impact on different phases of the BS procedure. The parameters are the number of nodes allowed at a given level (N), the number of greedy solutions for each node (q) and the maximum number of children per node (w).
In the context of the proposed BS algorithm, an increase in the number of evaluated nodes, driven by a deeper or wider search tree, directly impacts CPU time. Consequently, the initial parameter to consider is $N$, which represents the number of nodes allowed within a specified level. The challenge lies in selecting a value that ensures an adequate array of potential initial solutions while keeping the tree’s size within reasonable limits, which would otherwise result in higher CPU time. This effect becomes even more critical as the problem instance increases in complexity, particularly with the addition of more ports, vessels, or a longer planning horizon. Although a more extensive exploration of the solution space may be acceptable for small instances, it can quickly become computationally prohibitive for larger ones. Therefore, careful calibration of the tree structure is essential to ensure a proper balance between solution quality and computational efficiency.
The Successor Evaluation phase follows the Node Expansion phase. Comparing partial solutions in the MIRP is a high challenge. Therefore, during the evaluation of successors, different criteria were tested. Comparison of incomplete solutions can lead to imprecise and unsatisfactory evaluations. As a result, a greedy heuristic is applied to evaluate the quality of each generated successor. Thus, the second parameter, denoted as $q$, specifies the number of greedy solutions generated during the process. Specifically, the greedy heuristic is executed once following a deterministic approach, and subsequently, it is repeated $q-1$ times incorporating randomness into its execution to introduce variability in the solution space. To select the best successors, the median of the associated $q$ values is used as a representative measure. The third parameter, denoted as $w$, represents the upper limit on the number of children that can be generated by a single node. This parameter plays an important role in maintaining the diversity and heterogeneity among the solutions.
@@
One problem already mentioned in Section 4.1 is the fact that several identical solutions are represented in different ways. A simple procedure was devised with the aim of reducing the occurrence of solutions that share the same cost, thereby improving the efficiency of BS, as it is undesirable to occupy unnecessary space in the BS level. This procedure functions by keeping a list in which the keys correspond to the scores of potential solutions. Referencing this list, it is possible to identify whether a particular score has been assigned to any solution. In such cases, these solutions are excluded from further consideration. An unlikely issue may arise with the presence of two distinct solutions that share the same cost, due to the intricate nature of the problem. Nevertheless, even if this occurs, the benefits continue to outweigh the disadvantages.
@@ 4.4 The Greedy Algorithm
To effectively compare partial solutions, i.e., the successors produced during a BS execution, a greedy heuristic is used to construct a complete solution starting from a given partial solution, as presented in Algorithm 1. The heuristic builds a feasible solution using a greedy method, making each decision by selecting the most promising immediate choice according to a defined criterion. The criterion used involves choosing the next port at risk of violating inventory levels and identifying which vessel can arrive there first. Therefore, the path is designed to prioritize, in the subsequent call, the inclusion of the port at risk of inventory violation and the vessel capable of reaching this port in the earliest possible time.
@@
3while  $t\leq$ time_horizon  do
4       next_port $\leftarrow$ next port to violate
5      next_vessel $\leftarrow$ vessel arrives first
6      $s_{partial}\leftarrow s_{partial}\ \cup$ (next_port, next_vessel)
@@
To enhance flexibility and solution exploration, a nondeterministic version of the greedy algorithm is implemented, introducing randomness at two stages of the process, the port selection and the vessel selection. This randomness mechanism allows the algorithm to explore a wider range of possible paths, helping to diversify the solutions generated by the constructive method and potentially increasing the chances of finding better solutions compared to the deterministic version. Instead of choosing the port that first violates its capacity constraint or the ship that first serves the selected port, ports with earlier violation times and vessels with earlier arrival times have a higher probability of being selected.
By incorporating a probability factor within a normal distribution and considering the standard deviation, it is possible that the next port on the route might not be the first to face inventory violation, and the selected vessel may not be the one that reaches there first. The normal distribution was used to sample values around the average, introducing variation but ensuring that the majority of samples close to that remained near the central value.
Notably, throughout the BS procedure, multiple complete solutions are built using this greedy heuristic. To ensure that valuable solutions are kept, we keep the best $N$ solutions identified by the greedy heuristic, in case they outperform the final solution of the BS.
@@ 4.5 Local search
The local search procedure is based on a Randomized Variable Neighborhood Descent (RVND) scheme, where each neighborhood is explored in sequence until an improvement is found. If no improvement is found, the procedure moves to the next neighborhood, following a random order. If any neighborhood results in a better solution, the process is restarted. Furthermore, our implementation uses the first improvement strategy.
An important consideration when considering these neighborhoods is that in some cases certain calls may be displaced outside the planning horizon. When this issue occurs, these calls are ignored during the solution evaluation, as they do not contribute to the cost or feasibility within the planned period.
Six neighborhood structures are implemented, as described as follows.
@@
  • Remove: removes a call along with the next call of the same vessel. This procedure is carried out to preserve vessel parity. For example, if the call (P3, V1) is removed, it becomes necessary to remove the next call of V1 to maintain the producer-consumer balance.
@@ 4.6 Iterated Local Search
ILS (Lourenço et al., 2003) applies an iterative process that alternates between perturbation and local search procedures. In each iteration, a perturbation is randomly applied to the current solution, which involves making a controlled change in its structure, by altering the routing structure of the current solution, triggering a change in the cost value of the solution. The perturbation mechanism selects one of the six neighborhoods defined in Section 4.5 and applies a random move. In our implementation, such a mechanism is called twice in a row. This perturbation generates a new candidate solution which is then refined using a local search to improve its quality within the neighborhood of the current solution. The algorithm then checks if the new solution will be accepted by a pre-established Simulated Annealing criterion (Kirkpatrick et al., 1983). Once a solution is accepted, the algorithm updates the current solution and checks for improvement. If no improvement occurs, a counter increments. When this counter exceeds the allowed number of non-improving iterations, the solution is restored, and the counter resets. This process repeats until the maximum number of iterations is reached, at which point the final solution is returned.
@@ 5 Computational Experiments
The proposed algorithm was coded in Julia 1.10.5 and executed on an Intel(R) Core(TM) i7-8700K CPU @ 3.70GHz running Ubuntu 20.04 on a single thread.
@@ 5.1 Instance Details
Benchmark instances and their best-known solutions are maintained on the MIRPLib website [^1] http://mirplib.scl.gatech.edu. The metadata includes high-level information for each instance. Most of the parameters are self-explanatory and were previously presented in Section 3. The reader is referred to Papageorgiou et al. (2014) for further explanation. A software package has been created to read and use these public instances. Upon publication of the article, this package will be published.
This paper addresses Group 2 instances, representing a single-product MIRP, considering only one port per region and never involving split pickups or deliveries. All instances have one loading region (LR) and different numbers of discharging regions (DR). For example, instance LR1DR02VC01V6a has one loading region, two discharging regions, six vessels of one class, while instance LR1DR04VC03V15b has one loading region, four discharging regions, 15 vessels of three classes. The final letter “a” or “b” refers to the production/demand rates in the ports. If the letter is “a” the rates are constant, and if it is “b” the rates vary during the time horizon.
According to the classification presented by Papageorgiou et al. (2018), instances are grouped into three distinct categories based on their difficulty: Easy (E), Medium (M), and Hard (H). An instance is categorized as hard if, within a time limit of 1800 seconds, no commercial MILP solver operating with default parameter settings obtains a reduction greater than 10% of the gap. In contrast, an instance is considered easy if at least one commercially available MILP solver can reduce the gap by more than 90% within the same time limit using the default solver settings. Instances that do not meet the criteria for the easy or hard categories are classified as medium.
@@ 5.2.1 BS Parameters
To establish a promising value for the maximum number of children per node $w$, a set of experiments was carried out. The tests were carried out on easy instances, with $N=100$ and 10 seeds per instance. Table 1 reports the average and median objective value gaps obtained from tests in which the parameter $w$ was changed from 2 to 7. The percentages represent the aggregated gap in the objective values across all runs. The best performance was achieved with $w=2$, which resulted in the lowest average and median gaps.
Table 1: Gap analysis for different $w$ values.
@@ 5.2.2 Greedy Heuristic Parameters
Further tests were carried out to obtain the number of greedy solutions $q$ for each node, with the aim of maximizing performance relative to CPU time. To this end, the number of nodes allowed per level in the BS $N$ was set to be inversely proportional to $q$, ensuring that the CPU time remained approximately constant. Figure 2 illustrates the results, showing the calibration of $q$ for easy instances. The score indicates the average gap obtained across 10 different seeds. The best results were achieved for $q=3$, as it yielded the smallest average gap.
@@ 5.2.3 ILS Parameters
To tune the ILS parameters, the Julia HyperTuning.jl package (de Dios et al., 2021) was used for automated hyperparameter optimization. We selected 15 instances and five seeds, conducting a total of 120 trials that lasted 72 hours. The parameter range, step, and results are detailed in Table 2. The results indicated the final set of parameter values that corresponds to the configuration that achieved the best overall performance.
@@
Parameter Range Step Result
SA initial probability 0.75–0.99 $0.01$ $0.79$
SA final probability 0.01–0.10 $0.01$ $0.01$
ILS number of iterations 10–1000 10 640
ILS number of non-improving iterations before solution restore 1–20 1 4
ILS number of perturbations 1–5 1 2